# Cones of Noether–Lefschetz divisors and moduli spaces of hyperkähler manifolds

### v2

@@ 2.2. Rational Picard group of orthogonal modular varieties
A recently established key feature of our setting is that the $\mathbb{Q}$-vector space ${\rm{Pic}}_{\mathbb{Q}}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ can be seen as a space of vector-valued modular forms. This is what we explain now.
@@
The modular form $f$ is called a cusp form if $a_{0,\mu}=0$ for all isotropic elements $\mu\in D(\Lambda)$, i.e, the function $\sum_{m}a_{m,\mu}q^{m}$ vanishes at the cusp of $\mathbb{H}$. The function $f$ is called an almost cusp form if $a_{m,\mu}=0$ for all isotropic elements $\mu$ except possibly $0\in D(\Lambda)$ (see for instance [Pet15]*Section 3.3). Cusp forms and almost cusp forms form subspaces
@@
Let $\widetilde{\Gamma}_{\infty}$ be the stabilizer in ${\rm{Mp}}_{2}(\mathbb{Z})$ of the cusp at infinity. Assume further that $2k\equiv 2-n\mod 4$. Then for every half integer $k>2$ the Eisenstein series
@@
(6) $\varphi\colon\left({\rm{Mod}}_{k,\Lambda}^{\circ}\right)^{\vee}\longrightarrow{\rm{Pic}}_{\mathbb{Q}}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{{\rm{O}}}^{+}(\Lambda)\right),\;\;\;c_{m,\mu}\mapsto H_{-m,\mu}$
@@
Under the above isomorphism $\varphi$, the Hodge class $\lambda$ is identified with the functional $-c_{0,0}$ sending $E_{k,\Lambda}$ to $-1$ and $S_{k,\Lambda}$ to 00.
@@ 2.3. Effective and NL cones
It was shown in [BM19] that, on the left-hand side of (6), the functionals $c_{m,\mu}$ converge projectively to $-c_{0,0}$ as $m$ grows. This implies that the cone spanned by all $H_{m,\mu}$ is polyhedral. Using the formula (4), Bruinier–Möller moreover show that the cone ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ generated by primitive Heegner divisors $P_{\Delta,\delta}$ is polyhedral, answering [Pet15]*Question 4.5.2. More precisely, [BM19] shows that there is a neighborhood $\mathcal{U}$ of $\mathbb{Q}_{\geq 0}\lambda$ strictly contained in ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ and a value $\Delta_{0}$ such that for all $\Delta\geq\Delta_{0}$, we have $P_{\Delta,\delta}\in\mathcal{U}$. The NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ is then the convex hull of the divisors $P_{\Delta,\delta}$ for $\Delta\leq\Delta_{0}$.
@@
Confirming these formulas for a given $d$ requires explicitly computing the $\mathcal{U}$ and $\Delta_{0}$ described above. This has to do with finding concrete bounds analogous to Deligne’s bound for scalar-valued Hecke eigenforms of integral weight. Once these $\mathcal{U}$ and $m_{0}$ are computed, calculating ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ can be accomplished by computer. See Section 3 for more details.
@@ 3. Generators of the NL cone
As described in Section 2.3, in order to describe the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ for a given lattice $\Lambda$, one needs to calculate a neighborhood $\mathcal{U}$ of $\mathbb{Q}_{\geq 0}\lambda$ strictly contained in ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ and an explicit value $\Omega$ such that $P_{\Delta,\delta}\in\mathcal{U}$ for all $\Delta>\Omega$. Further, as in Subsection 2.2, we view ${\rm{Mod}}_{k,\Lambda}$ and $S_{k,\Lambda}$ as $\mathbb{Q}$-vector spaces.
@@
The coefficients of $e$ can be expressed in closed form [BK01] and a lower bound of the form $c_{m,\mu}(e)\geq C_{k,\Lambda}\cdot m^{k-1}$, where $C_{k,\Lambda}$ is an explicit positive constant depending only on the lattice $\Lambda$ and weight $k$, easily follows, cf. [BM19]*Propositions 3.2 and 4.5. As for cusp forms, despite the considerable literature on bounds for the growth of Fourier coefficients, we are unaware of a general bound (with explicit constants) that applies to our situation so we derive one below. The bound we derive is only the trivial bound $\mathrm{O}(m^{k/2})$, but this is sufficient to distinguish it from the growth of the lower bound for $c_{m,\mu}(e)$.
@@
$C=C(k)=\frac{(2\pi)^{k}}{\Gamma(k)\cdot(k-2)}+2.125.$
@@
$c_{m,\mu}(n,\beta)=2\pi\Big{(}\frac{m}{n}\Big{)}^{(1-k)/2}\sum_{c=1}^{\infty}\frac{1}{c}J_{k-1}(4\pi\sqrt{mn}/c)\cdot\mathrm{Re}\Big{[}e^{-\pi ik}K_{c}(\mu,m,\beta,n)\Big{]},$
where $K_{c}$ is the generalized Kloosterman sum
@@
(cf. [NIST:DLMF], 10.14.4). For small values of $c$ (say $c\leq n$), we use the first bound:
@@
$\left|c_{m,\mu}(n,\beta)\right|$ $\leq 2\pi\Big{(}\frac{m}{n}\Big{)}^{(1-k)/2}\cdot(4\pi)^{-1/3}m^{-1/6}M\cdot n^{7/6}+\frac{(2\pi)^{k}}{\Gamma(k)(k-2)}n.$
@@
$\left|c_{m,\mu}(m,\mu)\right|\leq 2^{1/3}\pi^{2/3}M\cdot m+\frac{(2\pi)^{k}}{\Gamma(k)(k-2)}\cdot m.$
The claim follows because $2.1252^{1/3}\pi^{2/3}M<$ 2.125. ∎
@@
We will first describe how to compute the cone of Heegner divisors ${\rm Eff}^{H}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}(\Lambda)\right)$. Let $\left({\rm{Mod}}_{k,\Lambda}^{\circ}\right)^{\vee}$ be the space of linear functionals on ${\rm{Mod}}_{k,\Lambda}^{\circ}$ and consider the cone $\mathcal{C}$ generated by the coefficient extraction functionals
@@
where $p(n)$ is the number of partitions of $n$. In particular the coefficient of $q^{m}$ in the expansion of $\Delta^{-b}$ is zero for $m<-b$ and the Fourier coefficients of the product $f(\tau)$ can be computed explicitly. We write
@@
As an example, we take the lattice $\Lambda=\Lambda_{4}$ corresponding to the moduli of degree four K3 surfaces. Then $S_{\frac{21}{2},\Lambda_{4}}$ is two dimensional generated by
@@
$\Delta^{-1}\cdot E_{\frac{7}{2},\Lambda(-1)}=q^{-1}\mathfrak{e}_{0}+64q^{-1/8}\mathfrak{e}_{\ell_{*}}+14q^{-1/2}\mathfrak{e}_{2\ell_{*}}+64q^{-1/8}\mathfrak{e}_{3\ell_{*}}+84\mathfrak{e}_{0}+\sum_{\begin{subarray}{c}(m,\mu)\\ m>0\end{subarray}}\alpha_{m,\mu}q^{m}\mathfrak{e}_{\mu}.$
@@
generates $S_{k,\Lambda_{4}}^{\vee}$. Then with $\lambda_{1}=\frac{1}{84},\lambda_{2}=\lambda_{4}=\frac{64}{84}$, and $\lambda_{3}=\frac{14}{84}$, Equation (9) holds.
We will need to transfer these results for the Petersson norm in terms of the $\ell^{2}$-norm on $\mathbb{Q}^{M}$. Recall that we identify each functional $s_{m,\mu}$ with the tuple
@@
Define an inner product on ${\rm{Mod}}_{k,\Lambda}^{\circ}$ as follows: for $f\in S_{k,\Lambda}$ then $\|f\|$ is the usual Petersson norm and we declare the Eisenstein series $E_{k,L}$ to have norm one and be orthogonal to $S_{k,\Lambda}$.
@@
These are convenient because their Petersson norm is easy to bound using Lemma 3.1. Indeed, writing $f_{m,\mu}=\sum c(n,\gamma)q^{n}\mathfrak{e}_{\gamma}$, one has
@@
Note that $\mathcal{C}_{S}$ contains an open neighborhood of 00 by [BM19]. To compute a concrete radius $R$, we write $\mathcal{C}_{S}\subset\mathbb{Q}^{M}$ as an intersection of finitely many half-planes, say $\{x:\,\langle v,x\rangle\leq a\}$, and take $R$ to be the minimum of $|a|/\|v\|_{\ell^{2}}$, where the latter is the standard $\ell^{2}$-norm on $\mathbb{Q}^{M}$. As for the choice of a constant $C_{k,\Lambda}$, when the discriminant of $\Lambda$ is $D$, it that can be derived from [BM19] is
@@
with the constant $B$ by (11). Recall that $\|e\|=1.$ Since $C_{k,L}$ is such that
@@
and $S_{k}\left(\Gamma_{1}(N^{2})\right)$ is the standard space of scalar-valued cusps forms of weight $k$ with trivial character for $\Gamma_{1}(N^{2})\subset{\rm{SL}}_{2}(\mathbb{Z})$.
@@
We now impose the added assumption that $\Lambda$ splits off two copies of the hyperbolic plane. Then the functionals $c_{m,\mu}$ correspond to the (non-primitive) Heegner divisors $H_{m,\mu}$ under the isomorphism (6) of Theorem 2.2. Hence in this case, Theorem 3.4 describes a generating set for the Heegner cone ${\rm Eff}^{H}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}(\Lambda)\right)$.
Continuing with the hypothesis that $\Lambda$ splits off two copies of the hyperbolic plane, we will now use the bounds of Theorem 3.4 in order to compute the NL cone ${\rm Eff}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}(\Lambda)\right)$.
To state the explicit bound $\Omega$ in the case of the $P_{\Delta,\delta}$ generating ${\rm Eff}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}(\Lambda)\right)$, define the functionals
$p_{\Delta,\delta}:=\sum_{\begin{subarray}{c}r\in\mathbb{Z}_{>0}\\ r^{2}|\Delta\end{subarray}}\mu(r)\sum_{\begin{subarray}{c}\sigma\in D(\Lambda)\\ r\sigma=\delta\end{subarray}}c_{\Delta/r^{2},\sigma},$
such that via the isomorphism (6) one has $\varphi(p_{\Delta,\delta})=P_{\Delta,\delta}$ is the corresponding primitive Heegner divisor by Equation (4). Let $\mathcal{P}$ be the cone generated by the $p_{\Delta,\delta}$. As in the case of the Heegner cone, using the isomorphism of Theorem 2.2, a description of the generators of $\mathcal{P}$ gives a description of the generators of ${\rm Eff}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}(\Lambda)\right)$.
@@
$(-7159053/4,-1/2),(-1/2,-3880799/602547),(2143005/2873041,122245370/979706981).$
@@
$R=\min\Big{(}\frac{3880800}{\|(2,602547)\|},\frac{13582800}{\|(38,-108856407)\|},\frac{2293200}{\|(-3175198,602547)\|}\Big{)}\approx 0.1248$
for the largest incircle centered at zero.
We have implemented Sage package [heegner_cones], which, given a lattice $\Lambda$ satisfying the given hypotheses of this section, applies method described above together with the bounds of Theorem 3.7 in order to compute the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$.
@@
We now explicitly compute ${\rm{Eff}}^{NL}\left(\mathcal{D}_{\Lambda}\big{/}\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ in some key examples.

### v3

@@ 2.2. Rational Picard group of orthogonal modular varieties
A recently established key feature of our setting is that the $\mathbb{Q}$-vector space ${\rm{Pic}}_{\mathbb{Q}}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ can be seen as a space of vector-valued modular forms. This is what we explain now.
@@
The modular form $f$ is called a cusp form if $a_{0,\mu}=0$ for all isotropic elements $\mu\in D(\Lambda)$, i.e, the function $\sum_{m}a_{m,\mu}q^{m}$ vanishes at the cusp of $\mathbb{H}$. The function $f$ is called an almost cusp form if $a_{m,\mu}=0$ for all isotropic elements $\mu$ except possibly $0\in D(\Lambda)$ (see for instance [Pet15]*Section 3.3). Letting ${\rm{S}}_{k,\Lambda}$ denote the space of cusp forms and ${\rm{Mod}}_{k,\Lambda}^{\circ}$ the space of almost cusp forms, we have subspace inclusions
@@
Let $\widetilde{\Gamma}_{\infty}$ be the stabilizer in ${\rm{Mp}}_{2}(\mathbb{Z})$ of the cusp at infinity. If $k>2$ is a half integer such that $2k\equiv 2-n\mod 4$, then the Eisenstein series
@@
(6) $\varphi\colon\left({\rm{Mod}}_{k,\Lambda}^{\circ}\right)^{\vee}\longrightarrow{\rm{Pic}}_{\mathbb{Q}}\left(\mathcal{D}\big/\widetilde{{\rm{O}}}^{+}(\Lambda)\right),\;\;\;c_{m,\mu}\mapsto H_{-m,\mu}$
@@
Under the above isomorphism $\varphi$, the Hodge class $\lambda$ is identified with the functional $c_{0,0}$ sending $E_{k,\Lambda}$ to 1-1 and $S_{k,\Lambda}$ to 0.
@@ 2.3. Effective and NL cones
It was shown in [BM19] that, on the left-hand side of (6), the functionals $c_{m,\mu}$ converge projectively to $c_{0,0}$ as $m$ grows. This implies that the cone spanned by all $H_{m,\mu}$ is polyhedral. Using the formula (4), Bruinier–Möller moreover show that the cone ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ generated by primitive Heegner divisors $P_{\Delta,\delta}$ is polyhedral, answering [Pet15]*Question 4.5.2. More precisely, [BM19] shows that there is a neighborhood $\mathcal{U}$ of $\mathbb{Q}_{\geq 0}\lambda$ strictly contained in ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ and a value $\Delta_{0}$ such that for all $\Delta\geq\Delta_{0}$, we have $P_{\Delta,\delta}\in\mathcal{U}$. The NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ is then the convex hull of the divisors $P_{\Delta,\delta}$ for $\Delta\leq\Delta_{0}$.
@@
Confirming these formulas for a given $d$ requires explicitly computing the $\mathcal{U}$ and $\Delta_{0}$ described above. This has to do with finding concrete bounds analogous to Deligne’s bound for scalar-valued Hecke eigenforms of integral weight. Once these $\mathcal{U}$ and $m_{0}$ are computed, calculating ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ can be accomplished by computer. See Section 3 for more details.
@@ 3. Generators of the NL cone
As described in Section 2.3, in order to describe the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ for a given lattice $\Lambda$, one needs to calculate a neighborhood $\mathcal{U}$ of $\mathbb{Q}_{\geq 0}\lambda$ strictly contained in ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ and an explicit value $\Omega$ such that $P_{\Delta,\delta}\in\mathcal{U}$ for all $\Delta>\Omega$. Further, as in Subsection 2.2, we view ${\rm{Mod}}_{k,\Lambda}$ and $S_{k,\Lambda}$ as $\mathbb{Q}$-vector spaces.
@@
The coefficients of $e$ can be expressed in closed form [BK01] and a lower bound of the form $c_{m,\mu}(e)\geq C_{k,\Lambda}\cdot m^{k-1}$, where $C_{k,\Lambda}$ is an explicit positive constant depending only on the lattice $\Lambda$ and weight $k$, easily follows, see [BM19]*Propositions 3.2 and 4.5. As for cusp forms, despite the considerable literature on bounds for the growth of Fourier coefficients, we are unaware of a general bound (with explicit constants) that applies to our situation so we derive one below. The bound we derive is only the trivial bound $\mathrm{O}(m^{k/2})$, but this is sufficient to distinguish it from the growth of the lower bound for $c_{m,\mu}(e)$.
@@
$C^{\prime}=C^{\prime}(k)=\frac{(2\pi)^{k}}{\Gamma(k)\cdot(k-2)}+2.125.$
For $m\geq 1$, we therefore have $|c_{m,\mu}(m,\mu)|\leq Cm$ with $C=C^{\prime}+1$. Lemma 3.1 then yields
$|a_{m,\mu}|\leq\tilde{C}m^{k/2}\cdot\|f\|.$
@@
$c_{m,\mu}(n,\beta)$ $=2\pi\Big(\frac{m}{n}\Big)^{(1-k)/2}\sum_{c=1}^{\infty}\frac{1}{c}J_{k-1}(4\pi\sqrt{mn}/c)\cdot\mathrm{Re}\Big[e^{-\pi ik}K_{c}(\mu,m,\beta,n)\Big]$
$+\frac{1}{2}\delta_{m=n,\mu=\beta}+\frac{1}{2}\delta_{m=n,\mu=-\beta},$
where $\delta$ is the Kronecker delta, $K_{c}$ is the generalized Kloosterman sum
@@
(see [NIST], 10.14.4). For small values of $c$ (say $c\leq n$), we use the first bound:
@@
$\left|c_{m,\mu}(n,\beta)\right|$ $\leq 2\pi\Big(\frac{m}{n}\Big)^{(1-k)/2}\cdot(4\pi)^{-1/3}m^{-1/6}M\cdot n^{7/6}+\frac{(2\pi)^{k}}{\Gamma(k)(k-2)}n+1.$
@@
$\left|c_{m,\mu}(m,\mu)\right|\leq 2^{1/3}\pi^{2/3}M\cdot m+\frac{(2\pi)^{k}}{\Gamma(k)(k-2)}\cdot m+1.$
The claim follows because $2^{1/3}\pi^{2/3}M<2.125$. ∎
@@
We will first describe how to compute the cone of Heegner divisors ${\rm Eff}^{H}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda)\right)$. Let $\left({\rm{Mod}}_{k,\Lambda}^{\circ}\right)^{\vee}$ be the space of linear functionals on ${\rm{Mod}}_{k,\Lambda}^{\circ}$ and consider the cone $\mathcal{C}$ generated by the coefficient extraction functionals
@@
where $p(n)$ is the number of partitions of $n$. In particular the coefficient of $q^{m}$ in the expansion of $\Delta^{-b}$ is zero for m<bm<-b and the Fourier coefficients of the product $f(\tau)$ can be computed explicitly. We write
@@
As an example, we take the lattice $\Lambda=\Lambda_{4}$ corresponding to the moduli of degree four K3 surfaces. Then $S_{\frac{21}{2},\Lambda_{4}}$ is two-dimensional generated by
@@
$\Delta^{-1}\cdot E_{\frac{7}{2},\Lambda(-1)}=q^{-1}\mathfrak{e}_{0}+64q^{-1/8}\mathfrak{e}_{\ell_{*}}+14q^{-1/2}\mathfrak{e}_{2\ell_{*}}+64q^{-1/8}\mathfrak{e}_{3\ell_{*}}+108\mathfrak{e}_{0}+\sum_{\begin{subarray}{c}(m,\mu)\\ m>0\end{subarray}}\alpha_{m,\mu}q^{m}\mathfrak{e}_{\mu}.$
@@
generates $S_{k,\Lambda_{4}}^{\vee}$. Then with $\lambda_{1}=\frac{1}{108},\lambda_{2}=\lambda_{4}=\frac{64}{108}$, and $\lambda_{3}=\frac{14}{108}$, Equation (9) holds.
We will need to rewrite these results for the Petersson norm in terms of the $\ell^{2}$-norm on $\mathbb{Q}^{M}$. Recall that we identify each functional $s_{m,\mu}$ with the tuple
@@
Define an inner product on ${\rm{Mod}}_{k,\Lambda}^{\circ}$ as follows: if $f\in S_{k,\Lambda}$ then $\|f\|$ is the usual Petersson norm and we declare the Eisenstein series $E_{k,L}$ to have norm one and be orthogonal to $S_{k,\Lambda}$.
@@
These are convenient because their Petersson norm is easy to bound using Lemma 3.1. Indeed, writing $f_{m,\mu}=\sum c(n,\gamma)q^{n}\mathfrak{e}_{\gamma}$, for any $m\geq 1$ one has
@@
Note that $\mathcal{C}_{S}$ contains an open neighborhood of 0 by [BM19]. To compute a concrete radius $R$, we write $\mathcal{C}_{S}\subset\mathbb{Q}^{M}$ as an intersection of finitely many half-planes, say $\{x:\,\langle v,x\rangle\leq a\}$, and take $R$ to be the minimum of $|a|/\|v\|_{\ell^{2}}$, where the latter is the standard $\ell^{2}$-norm on $\mathbb{Q}^{M}$. When the discriminant of $\Lambda$ is $D$, it can be derived from [BM19] that the constant $C_{k,\Lambda}$ can be chosen to be
@@
with the constant $B$ given by (12). Recall that $\|e\|=1.$ Since $C_{k,L}$ is such that
@@
and $S_{k}\left(\Gamma_{1}(N^{2})\right)$ is the standard space of scalar-valued cusp forms of weight $k$ with trivial character for $\Gamma_{1}(N^{2})\subset{\rm{SL}}_{2}(\mathbb{Z})$.
@@
We now impose the added assumption that $\Lambda$ splits off an additional copy of the hyperbolic plane. Then the functionals $c_{m,\mu}$ correspond to the (non-primitive) Heegner divisors $H_{m,\mu}$ under the isomorphism (6) of Theorem 2.2. Hence in this case, Theorem 3.4 describes a generating set for the Heegner cone ${\rm Eff}^{H}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda)\right)$.
Continuing with the hypothesis that $\Lambda$ splits off not just one but two copies of the hyperbolic plane, we will now use the bounds of Theorem 3.4 in order to compute the NL cone ${\rm Eff}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda)\right)$.
To state the explicit bound $\Omega$ in the case of the $P_{\Delta,\delta}$ generating ${\rm Eff}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda)\right)$, define the functionals
$p_{\Delta,\delta}:=\sum_{\begin{subarray}{c}r\in\mathbb{Z}_{>0}\\ r^{2}|\Delta\end{subarray}}\mu(r)\sum_{\begin{subarray}{c}\sigma\in D(\Lambda)\\ r\sigma=\delta\end{subarray}}c_{\Delta/r^{2},\sigma}.$
Then, applying the isomorphism (6), one has $\varphi(p_{\Delta,\delta})=P_{\Delta,\delta}$ is the corresponding primitive Heegner divisor by Equation (4). Let $\mathcal{P}$ be the cone generated by the $p_{\Delta,\delta}$. As in the case of the Heegner cone, using the isomorphism of Theorem 2.2, a description of the generators of $\mathcal{P}$ gives a description of the generators of ${\rm Eff}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}(\Lambda)\right)$.
@@
$\left(\frac{-7159053}{4},-\frac{1}{2}\right),\left(-\frac{1}{2},-\frac{3880799}{602547}\right),\left(\frac{2143005}{2873041},\frac{122245370}{979706981}\right).$
@@
$R=\min\left(\frac{3880800}{\|(2,602547)\|},\frac{13582800}{\|(38,-108856407)\|},\frac{2293200}{\|(-3175198,602547)\|}\right)\approx 0.1248$
for the largest circle centered at zero and contained in the given triangle.
We have implemented the Sage package [heegner_cones], which applies the method described above together with the bounds of Theorem 3.7 in order to compute the NL cone ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$.
@@
We now explicitly compute ${\rm{Eff}}^{NL}\left(\mathcal{D}\big/\widetilde{\rm{O}}^{+}\left(\Lambda\right)\right)$ in some key examples.